OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2) −18 18 −12 12

OpenStudy (anonymous):

Step 1. Distribute=> 4x +24=6x-12 Step 2. combine like terms=> 36=2x Step 3. Isolate x=> 18

OpenStudy (theviper):

4(x+6)=6(x-2) 4x+24=6x-12 6x-4x=24+12 2x=36 x=36/2=18 So the right answer is 18 :)

OpenStudy (theviper):

Understood??